OpenStudy (anonymous):

Choose the two equations you would use to solve the absolute value equation below. Then solve the two equations. |2x + 5| = 3 A. 2x + 5 = 3 and 2x + 5 = –3; {–1, –4} B. 2x + 5 = 3 and 2x + 5 = –3; {1, –4} C. 2x – 5 = 3 and 2x + 5 = –3; {–1, 4} D. 2x – 5 = 3 and 2x + 5 = –3; {–1, –4}

OpenStudy (anonymous):

When solving absolute value problems you have a "positive" version and a "negative" version. For example: for |2x+5| = 3 the positive is the equation as-is: 2x + 5 = 3, with a solution of x = -1. The negative is where the value of what's in the absolute value symbols would have been negative without those symbols. That is, for our example: -(2x + 5) = 3. This is the same as 2x + 5 = -3, which has a solution of x = -4. So you want choice A.

OpenStudy (anonymous):

ok thank you so much:)

OpenStudy (anonymous):

welcome :)